What is the equation of the line with a slope of -2/3 and passes through the point (-1, 4)
Thanks ?
And in what grade do they teach these stuff I want to know tnx

Answer by stanbon(75887)

What is the equation of the line with a slope of -2/3 and passes through the point (-1, 4)
------
Taught in 1st year Algebra.
-----
Equation Form: y = mx + b
---
You are told y = 4 when x = -1, and m = -2/3.
Solve for "b".
---
4 = (-2/3)(-1) + b
b = 4 - (2/3)
b = 10/3
----
Equation
y = (-2/3)x + (10/3)
==========================
cheers,
Stan H.